In the class of Statistics, there are 75 students in total out of which 55 male, 20 female students. 10 students need to be selected for the occasion of freshers' reception.



a. Determine the probability that 4 of the selected students would be girls by using the formula of binomial distribution.




1
Expert's answer
2022-06-07T15:32:39-0400

Let "X=" the number of girls: "X\sim Bin(n, p)."

Given "n=10, p=20\/75=4\/15, q=1-p=11\/15"



"P(X=4)=\dbinom{10}{4}(\dfrac{4}{15})^{4}(\dfrac{11}{15})^{10-4}""\approx0.16516"